A lottery system has balls numbered 1 to 55 and randomly selects 7 of the lottery balls. There is only one prize of $9,500,000.00$9,500,000.00which is awarded only if a lottery player selects the correct set of 7 lottery balls.

a) If a lottery ticket costs $5.00$5.00, what is a lottery player's expected value?

b) How much would the lottery prize have to be worth if it was to be a fair game?

(Note: Include dollar signs in your answer)
Answer #1

a) lottery player's expected value =expected win -cost of ticket

= 9500000*(1/55C7)-5 = 9500000*(1/202927725)-5=$ -4.95

b)

lottery prize for this to be fair =5*202927725=$1014638625